### v1

# Establishing Provenance Before Coding: Traditional and Next-Gen Signing

Abstract
Software engineers integrate third-party components into their applications. The resulting software supply chain is vulnerable. To reduce the attack surface, we can verify the origin of components (provenance) before adding them. Cryptographic signatures enable this. This article describes traditional signing, its challenges, and changes introduced by next generation signing platforms.
keywords:
Provenance; Software Signing; Software Re-Use; Component Selection; Cybersecurity; Software EngineeringNearly all modern software relies on other software components at some point [1]. As illustrated in Figure 1, these components include libraries, operating systems, and build and deployment tools. As a part of their planning process, software engineers decide which tools and dependencies they will use to build their software. Each of those components also has its own dependencies. A software supply chain forms as a result. Since these components can be developed by other teams or other organizations, the supply chain includes trust relationships between the product and its dependencies.
Engineers must consider functional (fit for purpose) and non-functional properties during component selection. One property of recent interest is cybersecurity, because the trust relationships within software supply chains have begun to be exploited by attackers. In recent years there have been many high-profile attacks and attempts [1]. As a result, many proposals have been made to enhance software supply chain security. In their literature review, Okafor et al.summarized those proposals as addressing three distinct properties: separation to ensure that failures in one component are isolated, transparency to see the full supply chain, and validity to prove that components are correct [2].
This article is focused on a class of techniques that increase transparency and validity: provenance. Among provenance techniques, signing stands out as the de facto method for assuring the origin of an artifact. Even though signing is an old concept, there have been many works showing that signing is easier said than done. However, next-generation signing platforms have emerged in recent years that improve usability.
In this article, we summarize software supply chains, the history of signing and traditional signing platforms, and the transformative aspects of next-generation software signing platforms. We discuss the current state of software signing and discuss how improvements to signing infrastructure are leading to better security practices. Our goal is for the reader to learn that, due to the changing software signing landscape, their component selection process can, and should, integrate provenance information.
“A few years from now, everybody expects provenance…they don’t install anything if they can’t prove where it came from…Signing for open source is probably the single most important thing for software supply chain security” — Technical leader111This and other quotes are taken from interviews of industry cybersecurity experts conducted by Kalu et al.in 2024 [6].
TRUST BEFORE CODING
Components Influence Product Code
Before writing code, software engineers consider the components they will use to build their software. One of the primary decisions is this: Which external components should we build on? Figure 1 depicts the kinds of components that software depends on, aligned with the Software Factory Model [4]. The first-party code in a software product must interact with external libraries — Which ones? It may rely on a framework or an operating system for interaction with the outside world — What is the resulting control flow, and which parts will be handled by the substrate? When being built — Should it make use of Ant or Gradle or Bazel? When being deployed — Should it rely on Docker, Podman, or Singularity as its virtualization scheme?
Each of these decisions affects the resulting first-party code. The selected libraries and substrates have an obvious effect: the code will call one API or another. More subtle influences are exercised downstream. Build tools may influence which language features are used, e.g., if the product can make use of GCC extensions to C or must rely on the C90 standard. Deployment tools may influence the parameterization of the system, e.g., the size of buffers and caches (and which of these knobs to expose).
Since dependency selection is a weighty matter, engineers evaluate candidate components along many dimensions. Functional considerations such as correctness and performance have long dominated this selection process. However, the rise of cyberattacks has made cybersecurity, and particularly the security risks of depending on external components, a top-of-mind concern.
Software Supply Chain Security
In the past, assessing the provenance of a component was left to the judgment of the engineering team. Now, the urgency of cyberattacks has driven the development of techniques to improve software supply chain security. To this end, security requirements related to component provenance are being introduced through industry standards and government regulations. Relevant industry standards include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [4], CNCF’s Software Supply Chain Best Practices, and Microsoft’s Secure Supply Chain Consumption Framework (S2C2F). The US government has also published standards such as NIST’s Secure Software Development Framework (NIST-SSDF) and NIST 800-204D, along with regulations for contractors handling government contracts (US Executive Order 14028 [5]).
As discussed earlier, Okafor et al.examined these proposals and distinguished three orthogonal properties that they seek to provide: separation, transparency, and validity [2]. An ideal software supply chain security strategy would satisfy all security properties. Most individual security techniques will promote only one of these properties. For example, containerization promotes separation between entities. Software Bill of Materials (SBOM) and software attestations ensure transparency. Software signing guarantees validity. In practice, a combination of techniques is often employed to strengthen the guarantee of one or more of these security properties. For example, signing build attestations (validity and transparency) provides downstream consumers with provenance.
In this article, we focus on a technique underpinning provenance — signing. Like many security practices, signing is not a silver bullet. Although signing does not guarantee the correctness of a piece of software, it does provide a way to verify the origin of a piece of software. Most industry standards also affirm the importance of signing, as it features heavily in their recommendations. For example, NIST 800-204D requires that attestations about software products (e.g., their bills of materials/SBOMs, build processes, etc.) be created by cryptographically signing with a secure key. NIST standards further require that consumers of software components verify these attestations before using them. Recently, several changes have been made to signing infrastructure that have improved the state of signing. This article focuses on the shortcomings of traditional signing methods and how newer techniques are attempting to resolve these issues.
Approaches to Establishing Provenance in Software Supply Chains
Several techniques exist to establish provenance in software supply chains. Examples include software bills of materials (SBOMs), attestations, and code signing. Many of these techniques use some form of signing to ensure the integrity of the information they provide. This enables end users to verify that information has not been tampered with.
Software Bills of Materials (SBOMs)
SBOMs contain a list of all the components that make up a piece of software. This list often includes the version of each component and the origin of the component. SBOMs can be used by downstream users to understand the risks associated with a piece of software. Furthermore, SBOMs can be used to ensure that a piece of software is built with the correct components. SBOMs can be signed to prove that the manifest has not changed.
Attestations
Attestations are pieces of metadata that describe some aspect of a software artifact. For examples, build attestations describe how a piece of software was built, and review attestations indicate that a particular person has reviewed a piece of code, a component, a process, etc. Typically, these attestations are signed to ensure their integrity. This enables users to verify that the metadata has not been tampered with.
Code Signing
Code signing is a method of verifying the origin of software artifacts. Users use digital signatures on source code, binaries, or even git commits to establish provenance. Downstream users can verify the signature to ensure that the software artifact came from an expected source. Although code signing does not make any quality guarantees about the software, it does provide a way to verify the origin of a piece of software.
TRADITIONAL SIGNING
Many platforms, such as software registries, support “traditional” signing as a way for engineers to verify the origin of the software components they rely on. Traditional signing is relatively simple but has drawbacks. When used correctly, traditional signing methods can cryptographically ensure the origin of an artifact. Unfortunately, these methods suffer from poor adoption and key management issues that can limit practical use.
How Traditional Signing Works
Several implementations of traditional signing exist. They share a common form. Figure 2 depicts the protocol for traditional software signing [10]. First, an author creates a package (A) and a public/private key pair (B). Next, the author uses the private key to sign the package (C). Then, the signed package (C) and the associated public key (D) are published. To use a package, an engineer downloads it (E) and its public key (F) and verifies the signature (G).
An Example: Pretty Good Privacy (PGP)
There are several established signing methods, such as Pretty Good Privacy (PGP), Git commit signing, and Docker Content Trust. These tools vary in the way that they implement the signing. Pretty Good Privacy (PGP), is a popular and well established format for traditional signing. This type of signing is implemented in a few different ways, but a notable instance is GNU Privacy Guard (GPG).
Using GPG, artifact authors can create a key with gpg --gen-key.
This walks them through creating a public/private key pair.
After generating a key pair, they can then sign some file with a command like gpg -ab some.file.
This creates a detached signature file called some.file.asc.
The author can then publish their file, signature, and public key.
After fetching the file, signature, and public key, an engineer downstream can verify the file using gpg --verify some.file.asc some.file.
Challenges
“A lot of the ways in which I think previous teams’ software signing hasn’t been useful is…key management…how the private key is managed but also in how the public key is made available” — Member of senior management
The example commands of GPG above seem simple. In just two commands, a software producer can create keys and a signature, and a consumer can validate the signature in a single command. However, there are several potential problems with this process.
Traditional signing methods have often been criticized for their lack of usability. Academic literature, including works similar to the “Johnny signs” studies [7], has documented these criticisms. In practice, there is a noticeable trend where several traditional signing tools are being discontinued by various software registries, such as PyPI’s cessation of PGP use [3]. Common usability issues are generally associated with key management, inadequate documentation, and user interface design challenges.
Key Management
Key management is particularly manual in traditional signing. Users are responsible for securely storing their private key and sharing their public key. This means they must not loose the private key and must keep it private. This becomes more complicated in organizational environments with personnel turnover. When keys become compromised, or are otherwise no longer secure, all corresponding signatures loose trust. This also necessitates the generation and distribution of new keys.
Identity Verification
Establishing trust in a public key is also particularly challenging. Once a user creates a key pair, they must find a way to share their public key. This is typically done through infrastructure like public key servers or through other file sharing means. Unfortunately, many of these methods do not provide strong identity guarantees for the creator of the public key. The web of trust was introduced as a way to introduce a more confident binding between an identity and a public key. There are, however, even issues with this technique (i.e., relying on social interactions like signing parties).
Transparency
Additionally, there are transparency issues with traditional signing schemes. Users typically do not have knowledge about which artifacts were signed or what information should be present. Without knowledge about what modifications have been made to a project’s metadata, users cannot verify that supply chain actors are behaving properly. For example, a downstream user may not be able to detect that the signature identity has changed in a project.
Status
The current state of signing varies depending on the type and environment of the software being developed. In general, signing is more prevalent in commercial software than in open source software.
Open Source
In open source, the state of signing is generally poor. This is depicted in Figure 3. The number of user created signatures on high profile platforms like Docker Hub, PyPI, and Hugging Face has historically been low. There are many potential reasons for this, but a simple explanation is that platforms do not require signatures [3]. This is particularly evident when comparing Maven Central (a platform that requires signing) to the rest in Figure 3.
Regardless of the reason, the small amount of signing on these platforms decreases the effectiveness of signing. Since so few people sign their projects, verifying signatures is very hard to do on a project relying on several dependencies. Furthermore, this lack of verification dissuades maintainers from going through the hassle of creating and managing signatures. PyPI is a prime example of this, after years of poor adoption, the registry maintainers decided to stop supporting PGP.
The case is a bit different in ecosystems that require signing. In instances like the Debian project or on platforms like Maven Central, contributors must sign and maintain keys. This means that other users of those ecosystems can verify provenance. Unfortunately, it also means that users of these platforms have to deal with the drawbacks of traditional signing schemes.
Commercial Software
In commercial software, the implementation and utilization of signing are significantly more pronounced. Recent surveys [9, 8] indicate that organizations value signing. However, the cost of setting up a signing infrastructure is a general concern. Although there is a diversity in the choice of tools, the perceived importance, and the challenges encountered, the adoption of signing as a practice remains widely recognized and actively employed.
NEXT-GENERATION SIGNING
“Sigstore comes in…keyless signing…you don’t have to worry about long-lasting static SSH keys, or keys being compromised, or rotating keys…you don’t have big servers to store all these signatures” — Manager
“[Sigstore’s] other strength…I can associate my identity with an OIDC identity as opposed to necessarily needing to generate a key and keep track of that key and yada, yada. So that’s super useful because I could say, ‘Oh, this is signed with [a] GitHub identity.’ So unless my GitHub identity has been compromised, that’s much better.”” — Member of senior management
In the past several years, many structural improvements have been made to signing infrastructure. We call the result next-generation signing tools.
Changes to Signing
Next-generation signing tools are motivated by the challenges associated with traditional signing methods, such as key management, identity verification, and transparency. These tools still use some of the primitive constructs (e.g., cryptographic algorithms) of traditional signing. However, they change the signing workflow depicted in Figure 2 to address shortcomings in traditional signing tools.
Key Management
One way to address key management issues is to use ephemeral keys. Ephemeral keys are keys that are generated on the fly and used for a single purpose. These keys do not need to be stored and managed by a user. This reduces the burden on users to ensure that they follow best practices for keys over their lifetime.
Identity Verification
Linking a public key to an identity is a difficult problem. Previous generation tools would verify identities on long-lived keys by validating key ownership off hand (e.g., by meeting with the keyholder in person during a conference). Next generation signing tools use new methods to bind an identity to an ephemeral key, and eventually to a signature. For example, OpenID Connect relies on the OAuth 2.0 protocol to verify the identity of a user. These methods tie a signature to an account, which can be used to verify the identity of the signer.
Shift Towards Key and Software Transparency
Transparency allows verifiers to ensure that key-issuers do not misbehave (e.g., by handing out a key to another party). Beyond allowing for easy key management for users, next generation signing systems integrate transparency for keys issued and signatures created. This allows users to actively monitor signatures and ensure that they are not being withheld, or re-used.
Examples of Next-Gen. Signing Platforms
Various systems have been engineered to fit the goals above in the last couple of years. Perhaps one of the most visible ones is the Sigstore project, which simplifies signature creation and verification. Sigstore is an open source and free to use project that is supported by the Linux Foundation and several other organizations. In order to achieve this, Sigstore leverages existing standards such as OIDC for identity verification, as well as Transparency Logs for software and key transparency.
Similarly to Sigstore, the Open PubKey (OPK) project provides a similar construction (iedentity-providers to issue single-use ephemeral keys). However, it tries to minimize infrastructure requirements. Though OPK is designed to be applicable to other usecases such as ssh authentication using identity providers and end-to-end encrypted chat, it is also applicable for next-generation software signing. Currently, OPK is used by DockerHub to sign their official images [14]. A generalization of this construction is also being developed by standards organizations. Of particular interest, the IETF’s Proof of Issuer Key Authority (PIKA), can be applied to the same usecases as Sigstore and OPK [13].
Status
Next-generation software signing solutions have experienced a significant rise in popularity since their introduction. Industry (e.g., companies like Shopify, Autodesk and Verizon) and Open Source (e.g., the Python interpreter and Kubernetes) alike have integrated this schemes [11, 12]. These successes have lead to public recognition from government strategies such as the US government CISA’s “Improving Security of Open Source Software in Operational Technology and Industrial Control Systems” as well as industry organizations such as JPMorgan’s CISO. However, concrete measurements of adoption of these technologies is lacking, and thus it is hard to fully assess whether all this visibility is affecting adoption.
As a step toward quantifying the adoption status of next-generation signing tools, we looked at trends related to one of these tools, Sigstore. Our goal was to estimate its usage. We examined two aspects of usage: download rates of the Sigstore toolkit, and analyzing use and discussion of the tools and associated techniques, such as npm’s provenance feature,222See https://docs.npmjs.com/generating-provenance-statements. which use Sigstore under the hood.
To estimate the adoption rate of next-generation signing tools, we obtained the download statistics for various Sigstore-related tooling over the past 24 months. Two Sigstore tools were checked: NPM and PyPI. The PyPI download data was collected using the PyPI database as queried through Google’s BigQuery. The NPM download data was collected using npm-stat. Figure 4 summarizes our result. Both PyPI and npm versions of the tooling are seeing increasing use, measured in the tens of thousands to millions of downloads over time.
To observe the usage and discussion of next generation signing tools, we used Sourcegraph’s Code Search to query GitHub repositories for keywords relating to the tools. These keywords were determined based on package names (top half of the table) and indicators of provenance workflows within repositories (bottom half of the table). The results of these keyword searches are shown in Table 1.
| Approach | Measure | Mentions | Repos |
| sigstore | Name | 143,830 | 5,229 |
| sigstore-python | Name | 1,041 | 396 |
| sigstore-java | Name | 194 | 21 |
| sigstore-rs | Name | 235 | 23 |
| in-toto | Name | 68,752 | 985 |
| npm provenance workflow | GitHub Workflow | 37,363 | 26,255 |
| npm repos publishing with provenance | GitHub Workflow | 1,427 | 896 |
[…]
CONCLUSIONS
Software supply chains are not new, but they are more visible than ever before. Software signing — of code, of attestations, of bills of materials, and of all other aspects of consuming a software component — is a key technique in ensuring that supply chain artifacts have not been modified since they were authored by trusted parties. Although the previous generation of signing tools left much to be desired, the next generation of tools has addressed many of those challenges and is seeing widespread adoption.
The technological capability is now present for any component provider to easily sign their products. Software signature creation and verification are already commonplace. While selecting components, software engineers can and should now rely only on components bearing the signature of the author. Conversely, software producers should expect that their consumers will begin to demand these signatures, as a result of due diligence or the need to meet current or anticipated regulations regarding software provenance. Soon, checking signatures before coding will be a normal part of component selection.
Finally, we remind the reader that provenance — even cryptographically-assured provenance using signatures — is not a guarantee of correctness nor security. Provenance information reduces the risk of certain classes of attacks, such as man-in-the-middle substitutions. Provenance information also improves the engineer’s ability to estimate the quality of the components they rely on, e.g., based on the reputation of the supplier. But guaranteeing the correctness and security of a component (not to mention the resulting system) is a task for formal methods, which is another endeavor for computing.
ACKNOWLEDGMENTS
The authors thank Zach Steindler of GitHub for his feedback on our estimates of npm provenance adoption. This work was supported by the US National Science Foundation under award #2229703, and by funds from Google and Cisco.